# On the importance of parallel magnetic-field fluctuations for electromagnetic instabilities in STEP

### v1

@@ Appendix B Cancellation of Magnetic Drift and $\mathbf{B_{\parallel}}$ Terms in Appropriate Limits
$T_{\nabla B}^{p^{\prime}}$ $=$ $\frac{1}{q_{s}B_{0}}\frac{\mu_{0}\mathbf{k}_{\perp}\cdot\mathbf{b}_{0}\times\nabla p}{B_{0}^{2}}\frac{m_{s}v_{\perp}^{2}}{2}g_{s}=\omega_{\rm dia}\frac{\mu_{0}n_{0s}T_{0s}}{B_{0}^{2}}\frac{m_{s}v_{\perp}^{2}}{2T_{0s}}g_{s}$ (24)
where $\omega_{\rm dia}=\mathbf{k}_{\perp}\cdot\mathbf{V_{dia}}$ and the diamagnetic velocity $\mathbf{V_{dia}}=\mathbf{b}_{0}\times\mathbf{\nabla}p/(n_{0s}q_{s}B_{0})$.
@@
$T_{ds}$ $=$ $-\frac{1}{\Omega_{s}}\left[\omega_{\nabla B}\left(v_{\parallel}^{2}+\frac{v_{\perp}^{2}}{2}\right)+\frac{\mu_{0}(\mathbf{k}_{\perp}\cdot\mathbf{b}_{0})\nabla p}{B_{0}^{2}}v_{\parallel}^{2}\right]g_{s}.$ (25)
@@
In the long wavelength high frequency limit ($b_{s}\rightarrow 0$ and $\omega\gg\omega_{\star s}^{T}$) the compressional magnetic perturbation term on the RHS of the GKE reduces to:
$T_{B_{\parallel}}$ $=$ $\omega\frac{\delta\!B_{\parallel}}{B_{0}}\frac{m_{s}v_{\perp}^{2}}{2T_{0s}}F_{0s}=-\omega\frac{\mu_{0}\delta\!P_{\perp}}{B_{0}^{2}}\frac{m_{s}v_{\perp}^{2}}{2T_{0s}}F_{0s},$ (27)
@@
Consideration of MHD-1: In the long wavelength fluid limit, $g_{s}$ is Maxwellian to leading order in $\omega_{ds}/\omega$, so $T_{\nabla B}^{p^{\prime}}$ and $T_{B_{\parallel}}$ have approximately the same structure in velocity space. A cancellation between $T_{\nabla B}^{p^{\prime}}$ and $T_{B_{\parallel}}$ will arise at leading order if:
@@
This approximate cancellation between $T_{\nabla B}^{p^{\prime}}$ and $T_{B_{\parallel}}$ allows the $\nabla p$ contribution to the $\nabla B$ drift and $B_{\parallel}$ to be dropped from the GKE, and corresponds to MHD-1.
Consideration of MHD-2: On the other hand, $T_{\kappa}^{\nabla p}$ and $T_{B_{\parallel}}$ have manifestly different structure in velocity space, so exact cancellation is not possible in the GKE. While MHD-2 implements 9 to allow $\delta\!B_{\parallel}$ to be dropped from the GKE, it only does this at the expense of changing the local equilibrium and its magnetic curvature, $\mathbf{b}_{0}\cdot\nabla\mathbf{b}_{0}$. In effect MHD-2 changes the local equilibrium field curvature, and again removes the pressure gradient contribution to the $\nabla B$ drift, $T_{\nabla B}^{p^{\prime}}$ defined in 24, for this modified local equilibrium!
@@ Appendix C A useful identity involving the magnetic-drift frequency
$\omega_{ds}$ $=\mathbf{k}_{\perp}\cdot\mathbf{b}_{0}\times\left(\frac{v_{\parallel}^{2}}{\Omega_{s}}(\mathbf{b}\cdot\nabla)\mathbf{b}+\frac{v_{\perp}^{2}}{2\Omega_{s}}\nabla\ln B\right)=\frac{\omega_{\kappa}v_{\parallel}^{2}}{\Omega_{s}}+\frac{\omega_{\nabla B}v_{\perp}^{2}}{2\Omega_{s}}$ (29)
Using Ampère’s law, we can write
@@
where $\mathbf{b}$ is a unit vector aligned to the magnetic field, so that $\mathbf{b}\times(\nabla\times\mathbf{b})=-\mathbf{b}\cdot\nabla\mathbf{b}$. Now invoking equilibrium force balance $\mathbf{J}\times\mathbf{B}=\nabla p$ [see equation (128) of [21]] in (30) gives:
@@ Appendix D Snapshots of the turbulent fields
Figure 10: Snapshot contour plots at the outboard midplane ($\theta=0$) from the final time step of different nonlinear simulations discussed throughout the text: (a, d, g) $e\delta\!\phi/(\rho_{\star}T_{e}),$ (b, e, h) $\delta\!A_{\parallel}/(\rho_{\star}\rho_{s}B_{0})$, and (c, f, i) $\delta\!B_{\parallel}/(\rho_{\star}T_{e}B_{0}).$ Each field is normalised to the maximum value of $e\delta\!\phi/(\rho_{\star}T_{e})$ at the outboard midplane.

### v2

@@ Appendix B Cancellation of Magnetic Drift and $\mathbf{\delta\!B_{\parallel}}$ Terms in Appropriate Limits
$T_{\nabla B}^{p^{\prime}}$ $=$ $\frac{1}{q_{s}B}\frac{\mu_{0}\mathbf{k}_{\perp}\cdot\mathbf{b}\times\nabla p}{B^{2}}\frac{m_{s}v_{\perp}^{2}}{2}g_{s}=\omega_{\rm dia}\frac{\mu_{0}n_{0s}T_{0s}}{B^{2}}\frac{m_{s}v_{\perp}^{2}}{2T_{0s}}g_{s}$ (25)
where $\omega_{\rm dia}=\mathbf{k}_{\perp}\cdot\mathbf{V_{dia}}$ and the diamagnetic velocity $\mathbf{V_{dia}}=\mathbf{b}\times\mathbf{\nabla}p/(n_{0s}q_{s}B)$.
@@
$T_{ds}$ $=$ $-\frac{1}{\Omega_{s}}\left[\omega_{\nabla B}\left(v_{\parallel}^{2}+\frac{v_{\perp}^{2}}{2}\right)+\frac{\mu_{0}(\mathbf{k}_{\perp}\cdot\mathbf{b}\times\nabla p)}{B^{2}}v_{\parallel}^{2}\right]g_{s}.$ (26)
@@
In the long wavelength high frequency limit ($b_{s}\rightarrow 0$ and $\omega\gg\omega_{\star s}^{T}$) the compressional magnetic perturbation term on the RHS of the GK equation reduces to:
$T_{\delta\!B_{\parallel}}$ $=$ $\omega\frac{\delta\!B_{\parallel}}{B}\frac{m_{s}v_{\perp}^{2}}{2T_{0s}}F_{0s}=-\omega\frac{\mu_{0}\delta\!P_{\perp}}{B^{2}}\frac{m_{s}v_{\perp}^{2}}{2T_{0s}}F_{0s},$ (28)
@@
Consideration of MHD-1: In the long wavelength fluid limit, $g_{s}$ is Maxwellian to leading order in $\omega_{ds}/\omega$, so $T_{\nabla B}^{p^{\prime}}$ and $T_{\delta\!B_{\parallel}}$ have approximately the same structure in velocity space. A cancellation between $T_{\nabla B}^{p^{\prime}}$ and $T_{\delta\!B_{\parallel}}$ will arise at leading order if:
@@
This approximate cancellation between $T_{\nabla B}^{p^{\prime}}$ and $T_{\delta\!B_{\parallel}}$ allows the $\nabla p$ contribution to the $\nabla B$ drift and $\delta\!B_{\parallel}$ to be dropped from the GK equation, and corresponds to MHD-1.
Consideration of MHD-2: On the other hand, $T_{\kappa}^{\nabla p}$ and $T_{\delta\!B_{\parallel}}$ have manifestly different structure in velocity space, so exact cancellation is not possible in the GK equation. While MHD-2 implements 9 to allow $\delta\!B_{\parallel}$ to be dropped from the GK equation, it only does this at the expense of changing the local equilibrium and its magnetic curvature, $\mathbf{b}\cdot\nabla\mathbf{b}$. In effect MHD-2 changes the local equilibrium field curvature, and again removes the pressure gradient contribution to the $\nabla B$ drift, $T_{\nabla B}^{p^{\prime}}$ defined in 25, for this modified local equilibrium!
@@ Appendix C A useful identity involving the magnetic-drift frequency
$\omega_{ds}$ $=\mathbf{k}_{\perp}\cdot\mathbf{b}\times\left(\frac{v_{\parallel}^{2}}{\Omega_{s}}(\mathbf{b}\cdot\nabla)\mathbf{b}+\frac{v_{\perp}^{2}}{2\Omega_{s}}\nabla\ln B\right)=\frac{\omega_{\kappa}v_{\parallel}^{2}}{\Omega_{s}}+\frac{\omega_{\nabla B}v_{\perp}^{2}}{2\Omega_{s}}$ (30)
Using Ampère’s law for the equilibrium, we can write
@@
where $\mathbf{b}$ is a unit vector aligned to the equilibrium magnetic field, so that $\mathbf{b}\times(\nabla\times\mathbf{b})=-\mathbf{b}\cdot\nabla\mathbf{b}$. Now invoking equilibrium force balance $\mathbf{J}\times\mathbf{B}=\nabla p$ in (31) gives:
@@ Appendix D Snapshots of the turbulent fields
Figure 8: Snapshot contour plots at the outboard midplane ($\theta=0$) from the final time step of different nonlinear simulations discussed throughout the text: (a, d, g) $e\delta\!\phi/(\rho_{\star}T_{e}),$ (b, e, h) $\delta\!A_{\parallel}/(\rho_{\star}\rho_{s}B)$, and (c, f, i) $\delta\!B_{\parallel}/(\rho_{\star}T_{e}B).$ Each field is normalised to the maximum value of $e\delta\!\phi/(\rho_{\star}T_{e})$ at the outboard midplane. Simulations which return very large heat flux are typically associated with turbulent fields that exhibit radially-extended structures and reach large values of $(\delta\!A_{\parallel}T_{e})/(\delta\!\phi e\rho_{s}B).$